OpenStudy (anonymous):

how would u do this problem 2+square root 19-2x

zepdrix (zepdrix):

\[\large 2+\sqrt{19-2x}=?\]Is this what the problem looks like? Solve for x? Or just simplify? What is on the other side of the equation? :O

OpenStudy (anonymous):

yes that is the problem

OpenStudy (anonymous):

solve for x

zepdrix (zepdrix):

=0?

OpenStudy (anonymous):

your just bsolving for x

zepdrix (zepdrix):

you can't solve for x if we don't have an equality sign somewhere... :o

OpenStudy (anonymous):

it says x= that...would the problem just be undefined?

zepdrix (zepdrix):

\[\large 2+\sqrt{19-2x}=x\]Ok that's what we needed to know.... So this is what it should look like?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Hmm this one is a bit tricky :) We'll have to do a bunch of algebra to solve for x. First, subtracting 2 from each side gives us,\[\large \sqrt{19-2x}=x-2\]Then we'll square both sides to get rid of that square root,\[\large 19-2x=(x-2)^2\]

zepdrix (zepdrix):

Then you need to expand out that binomial :D Remember how to do those?\[\large (x-2)^2\qquad =\qquad (x-2)(x-2)\qquad =\qquad x^2-2x-2x+4\]

OpenStudy (anonymous):

so the final answer is x=-3 and x=5 correct?

OpenStudy (anonymous):

(-3,0) and (5,0)

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Ah yes very good! :) Sorry I had to go earlier D':